FriskIsNice:

Help? I have 2 more things to do and it’s 6:45 qwq I don’t wanna stay awake until 12 am qwq

FriskIsNice:

1 attachment
FriskIsNice:

*also I didn’t spam my internet is bugging ;-;*

FriskIsNice:

*cause sometimes people report me a using of spamming when I never mean to spam 😒*

anthonybuddy:

ok its alright

Angle:

what does n look like it is equal to?

Angle:

what does q look like it's equal to?

FriskIsNice:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Angle what does q look like it's equal to? \(\color{#0cbb34}{\text{End of Quote}}\) Well since it’s both negative it would equal a positive

Angle:

i- was talking about the number

FriskIsNice:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Angle i- was talking about the number \(\color{#0cbb34}{\text{End of Quote}}\) Flip but I was speaking the truth 😂

Angle:

double flip you were talking about the multiplication rule it doesn't work for addition or subtraction

FriskIsNice:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Angle double flip you were talking about the multiplication rule it doesn't work for addition or subtraction \(\color{#0cbb34}{\text{End of Quote}}\) Well heck- ALSO I WASNT CURSING I ANWSERED A WUESTION FOR ART ONCE AND I SAID FLIP AND SOMEONE SAID I FLIPPING CURSED 🤣

Angle:

so what does n look like it's equal to? (the number this time)

FriskIsNice:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Angle so what does n look like it's equal to? (the number this time) \(\color{#0cbb34}{\text{End of Quote}}\) I think -N = -4 and -q = -1?

Angle:

yes n = -4 and q = -1 in that case n - q = (-4) - (-1) = ?

FriskIsNice:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Angle yes n = -4 and q = -1 in that case n - q = (-4) - (-1) = ? \(\color{#0cbb34}{\text{End of Quote}}\) -3? Or do you gotta add by the opposite-

Angle:

-3 is right :) how about q - n = (-1) - (-4) = ?

FriskIsNice:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Angle -3 is right :) how about q - n = (-1) - (-4) = ? \(\color{#0cbb34}{\text{End of Quote}}\) I think positive 3

Angle:

yeee so we have n - q = -3 q = -1 q - n = 3

Angle:

now that should be enough info to answer your question

FriskIsNice:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @boredfr \(\color{#0cbb34}{\text{Originally Posted by}}\) @FriskIsNice \(\color{#0cbb34}{\text{Originally Posted by}}\) @Angle -3 is right :) how about q - n = (-1) - (-4) = ? \(\color{#0cbb34}{\text{End of Quote}}\) I think positive 3 \(\color{#0cbb34}{\text{End of Quote}}\) Quoting Angle : you were talking about the multiplication rule it doesn't work for addition or subtraction \(\color{#0cbb34}{\text{End of Quote}}\) *inhale* If your multiplying a negative number that’s like -1 - -4 i\(\color{#0cbb34}{\text{Originally Posted by}}\) @Angle now that should be enough info to answer your question \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{Originally Posted by}}\) @Angle now that should be enough info to answer your question \(\color{#0cbb34}{\text{End of Quote}}\)well my dears it was right 😂